OpenStudy (anonymous):

Winning percentage. In a basketball league, the Falcons won 15 of their first 20 games. In order to win 60% of the total number of games, how many more games will they have to play, assuming they win only half of the remaining games?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how

OpenStudy (anonymous):

15/20=75% winning 12/20=60% winning

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

well 15 of twenty thats .75 winning percentage an a record of 15-5 and a 60 percent is 5:3 so add one to each until you reach a ratio of 5:3 since all multiples of 5 end in 5 or zero just check wins of 5 10 15 20 15 30 etc. 20 10 25 15 5:3 answer is 20 more games??

OpenStudy (anonymous):

15/20 how does that determine the 75% and 12/20 60%?

OpenStudy (anonymous):

oooooh. im sorry, i misunderstood the question. just a second

OpenStudy (anonymous):

ok read over pls

OpenStudy (anonymous):

they will need to play 6 more games. this is because each game is worth 5% winning/losing, so to lower their percentage from 75% to 60, they need to lose 3 games, and because they win 50% of there games that would mean they have to play an additional 6 games to lower it to 60

OpenStudy (anonymous):

so the fraction would be 15/20 multiply the top and bottom by 5 to get a number out of 100 ( a percent ) so that would be 75/100. so if they win only HALF of the remaining games, 50/100 (50%) , we would have to add 10% more, so first we need a fraction as simplified form, 1/2, now the 10%, 10/100, so 1/10. now add 1/10 and 1/2. 1/2 multiply by 5 in top and bottom, 5/10, so plus 1/10 is 6/10. and since you need it out of 20 multiply by 2, so 6/10 x 2 = 12/20. So The Answer Is 12/20.

OpenStudy (anonymous):

so is this answer also incorrect, i may go with urs because it seems to make more sense

OpenStudy (anonymous):

i havent got a calculator on me to check my work, but my method should be correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

once you are satisfied with an answer you can press the 'good answer' button

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

ye, rubin is right

OpenStudy (anonymous):

ncorrect. Let's let x be the total number of games that need to be played. Then what expression can you use to represent how many remaining games there would be? Then, we want 75% of the 20 + 50% of the remaining games = 60% of the total. Can you try setting this one up again?